OpenStudy (anonymous):

S=2(lw+hw+hl) solve for h

OpenStudy (anonymous):

\[S = 2(lw+hw+hl)\]Divide both sides by 2: \[\frac{ S }{ 2 } = lw+hw+hl\]Now subtract the quantity lw: \[\frac{ S }{ 2 } - lw = hw + hl\]Now use distributive property to isolate h: \[\frac{ S }{ 2 } - lw = h(w+l)\]Now divide by quantity (w+l) for answer: \[\frac{ S }{ 2(w+l) } - \frac{ lw }{ (w+l) } = h\]If you want let's simplify to one fraction: \[\frac{ S-2lw }{ 2(w+l) } = h\] Hope this helps! :D